# Your Image is Secretly the Last Frame of a Pseudo Video

### v1

@@ 1 Introduction
While diffusion models can be viewed as a special case of a traditional sequential generative model, i.e., hierarchical variational autoencoders (HVAEs) (Sønderby et al., 2016a; Maaløe et al., 2019; Vahdat & Kautz, 2020), they tend to outperform standard HVAEs significantly in practice. The major differences between diffusion models and standard HVAEs are two-fold. First, diffusion models tend to have much more intermediate states, which may help improve the generation quality (Huang et al., 2021). Second, diffusion models incorporate exact self-supervised information for their intermediate states: they are supposed to match the corrupted (e.g., noisy or blurred) versions of the original target image at different corruption levels. These additional information helps regularize training and guide generation in diffusion models. On the other hand, the intermediate states in standard HVAEs are unobserved and one does not have explicit control of them. Consequently, there may be many different distributions over the intermediate states that are capable of generating images (i.e., the issue of unidentifiability (Locatello et al., 2019; Khemakhem et al., 2020)). The lack of identifiability of the intermediate states may poses challenges to the optimization during training since it suggests a huge hypothesis space with many sub-optimal solutions.
@@ 2.1 Preliminaries: Sequential Generative Models
where the factors $q_{\phi}(x_{t}|x_{t-1})$ and $q_{\phi}(x_{t-1}|x_{t},x_{0})$ are mean-field Gaussian distributions with mean and diagnoal variance parameterized by neural networks. Notably, HVAEs only require the prior over the last latent variable be a fixed distribution (e.g., standard Gaussian) as shown in Figure 1 and let the model figure out all the intermediate latent variables by maximizing the the tractable evidence lower bound (ELBO) with respect to the parameters of both generation and inference models:
@@ 2.2 Diffusion Model vs Hierarchical VAE
where $F_{h}$ is the matrix for simulating the heat equation until time $t$. This can be seen as an HVAE trained with explicit supervision signals from pseudo videos created by the heat equation. We create $T=18$ frames of pseudo videos for each training image. Figure LABEL:fig:heat demonstrates that the HVAE trained with pseudo videos created by the heat equation can generate much sharper and diverse digits than a standard HVAE which uses the same architecture but with a learnable encoder. Here, we deliberately use heat equation instead of Gaussian noise to create pseudo frames to show that there are plenty of choices to create pseudo videos that contain useful self-supervised information besides adding Gaussian noise as in standard diffusion models.
@@ 3 Improved Reconstruction and Generation in VQ-VAE with Pseudo Videos
In this section, we utilize pseudo videos to improve image generation quality of Vector Quantized Variational Autoencoder (VQ-VAE) (Van Den Oord et al., 2017), where the latent variables $z$ are discrete tokens. Specifically, we employ its video generative model counterpart C-ViViT (Villegas et al., 2022) to compress the pseudo videos into latent discrete tokens. C-ViViT is trained by reconstructing the pseudo videos with L2 reconstruction loss (Kingma & Welling, 2013), Vector Quantization (VQ) loss (Van Den Oord et al., 2017), GAN style adversarial loss (Karras et al., 2020), and image perceptual loss (Johnson et al., 2016; Zhang et al., 2018). For the latent space created by a C-ViViT, we consider two generative models to fit a prior $p(z)$ for the latent tokens:
@@
Network Architectures. [^1] For 1-frame models, we have tried using deeper architectures but observed no improvement in performance, which suggests that our pseudo video framework can help further improve the performance of generative models while simply increasing the model size becomes ineffective. We train VQ-VAE based generative models with a codebook size of 1024 for the discrete latent tokens. Specifically, we use C-ViViT as compression model (autoencoder) for pseudo videos. For a pseudo video with shape $(T,H,W,C)$, we compress it to discrete latent tokens with shape $(\frac{T}{2},\frac{H}{4},\frac{W}{4},C)$. For images with shape $(1,H,W,C)$, we use VQ-VAE to compress them to tokens with shape $(1,\frac{H}{4},\frac{W}{4},C)$. We then consider two video generative models, VideoGPT and Phenaki (and their image generative model counterparts, ImageGPT and MaskGit), to fit the prior over the latent tokens.
  • VQ-VAE/C-ViViT (reconstruction). We use a similar architecture as in Villegas et al. (2022), which has a 4-layer spatial Transformer and a 4-layer temporal Transformer, with a hidden dimension of 512. For 1-frame VQ-VAE model, we consider a 8-layer spatial Transformer.
@@
Results. Table 1 shows the last-frame FID of C-ViViT for reconstruction and that of VideoGPT and Phenaki for AR and masked generation on CIFAR10 and CelebA, respectively. The 1-frame results correspond to the performance of their image model counterparts (i.e., VQ-VAE for image reconstruction, and ImageGPT and MaskGit for AR and masked image generation, respectively). We observe that pseudo videos indeed help improve the training of the C-ViViT as the reconstruction quality of the last frame is significantly improved with a few more frames. We show reconstructed CIFAR10 and CelebA images from different C-ViViT models in Figures LABEL:fig:recon_cifar10 and LABEL:fig:recon_celeba, respectively. It can be seen that for both datasets, the reconstructed images trained with 1-frame models are over-smoothed and this issue is resolved by using pseudo videos which produce much sharper images. Quantitatively, reconstruction FID improves as more frames are used. Pseudo videos also improve the image generation performance compared to the 1-frame results. Interestingly, we see a diminishing return as we include more frames. For CelebA images, 18-frame pseudo videos help achieve the best image generation performance for both AR and latent masked generation, but 8-frame models achieve a comparable performance. For CIFAR10 images, 8-frame pseudo videos result in better image generation performance than 18-frame pseudo videos, which suggests the latent codes have a more complex prior distribution in order to reconstruct model pseudo videos with 18 frames well and therefore this prior is more difficult for VideoGPT or Phenaki to capture. In summary. while pseudo videos help improve the performance, the optimal number of frames may depend on the dataset and the augmentation strategy. This diminishing return is not a severe issue in practice since practitioners may prefer to improve the generation with just a few more pseudo frames to avoid introducing a high computational cost.
Table 1: Last-frame FID of images produced by C-ViViT (reconstruction), VideoGPT (AR generation) and Phenaki (latent masked generation) trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQ-VAE (reconstruction), ImageGPT (AR generation) and MaskGit (latent masked generation) trained on original CIFAR10 and CelebA images.
@@ 4.1 Is First-order Markov Chain the Optimal Choice for Creating Pseudo Videos?
Moreover, the non-optimality of creating pseudo video via first-order Markov chain becomes clear: the first-order Markov data augmentation implies that $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$ and consequently $\mathcal{L}_{2}^{*}=\mathcal{L}_{1}^{*}$. More specifically, for strict inequality in Eq 14, we need to avoid $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$, which is equivalent to avoiding the use of either first-order Markov chain $x_{T}\rightarrow x_{T-1}\rightarrow x_{T-2}$ or $x_{T}\leftarrow x_{T-1}\rightarrow x_{T-2}$. This analysis is informative for us to design better pseudo videos, for example through data augmentation with higher-order Markov chains. We formalize the above informal reasoning into Theorem 4.1 and provide the formal proof in Appendix A.
@@ 4.2 Experiments
  • First-order Markov. We add Gaussian noise recursively 3 or 7 times to create first-order Markov pseudo videos with a linear schedule (Ho et al., 2020) with $\beta$ ranging from 0.0001 to 0.05.
  • High-order Markov. While using the same noise schedule to create $x_{T-t}$, instead of adding Gaussian noise to $x_{T-t+1}$, we use a simple strategy to create high-order Markov pseudo videos by adding Gaussian noise to the mean of $\{x_{T-s+1}\}_{s=1}^{t}$.
@@ 6 Conclusion and Discussions
Appendix A Proof of Theorem 4.1
@@
Appendix B Reproducibility Statement

### v2

@@ 1 Introduction
While diffusion models can be viewed as a special case of a traditional sequential generative model, i.e., hierarchical variational autoencoders (HVAEs) (Sønderby et al., 2016a; Maaløe et al., 2019; Vahdat & Kautz, 2020), they tend to outperform standard HVAEs significantly in practice. The major differences between diffusion models and standard HVAEs are two-fold. First, diffusion models tend to have much more intermediate states, which may help improve the generation quality (Huang et al., 2021). Second, diffusion models incorporate exact self-supervised information for their intermediate states: they are supposed to match the corrupted (e.g., noisy or blurred) versions of the original target image at different corruption levels. These additional information helps regularize training and guide generation in diffusion models. On the other hand, the intermediate states in standard HVAEs are unobserved and one does not have explicit control of them. Consequently, there may be many different distributions over the intermediate states that are capable of generating images (i.e., the issue of unidentifiability (Locatello et al., 2019; Khemakhem et al., 2020)). The lack of identifiability of the intermediate states may pose challenges to the optimization during training since it suggests a huge hypothesis space with many sub-optimal solutions.
@@ 2.1 Preliminaries: Sequential Generative Models
where the factors $q_{\phi}(x_{t}|x_{t-1})$ and $q_{\phi}(x_{t-1}|x_{t},x_{0})$ are mean-field Gaussian distributions with mean and diagnoal variance parameterized by neural networks. Notably, HVAEs only require the prior over the last latent variable be a fixed distribution (e.g., standard Gaussian) as shown in Figure 1 and let the model figure out all the intermediate latent variables by maximizing the tractable evidence lower bound (ELBO) with respect to the parameters of both generation and inference models:
@@ 2.2 Diffusion Model vs Hierarchical VAE
Table 1: Inception Score (IS) of generated digits from HVAE with encoder fixed according to the heat equation and standard HVAE with learnable encoder.
HVAE with heat equation encoder Standard HVAE with learnable encoder
IS 9.32 7.27
@@
where $F_{h}$ is the matrix for simulating the heat equation until time $t$. This can be seen as an HVAE trained with explicit supervision signals from pseudo videos created by the heat equation. We create $T=18$ frames of pseudo videos for each training image. Figure LABEL:fig:heat demonstrates that the HVAE trained with pseudo videos created by the heat equation can generate much sharper and diverse digits than a standard HVAE which uses the same architecture but with a learnable encoder. We also report in Table 1 the Inception Score (IS) of the generated images for quantitative comparison and HVAE trained with pseudo videos created by the heat equation achieves noticeably higher IS than standard HVAE. Here, we deliberately use heat equation instead of Gaussian noise to create pseudo frames to show that there are plenty of choices to create pseudo videos that contain useful self-supervised information besides adding Gaussian noise as in standard diffusion models.
@@ 3 Improved Reconstruction and Generation in VQ-VAE with Pseudo Videos
In this section, we utilize pseudo videos to improve image generation quality of Vector Quantized Variational Autoencoder (VQVAE) (Van Den Oord et al., 2017), where the latent variables $z$ are discrete tokens. Specifically, we employ its video generative model counterpart C-ViViT (Villegas et al., 2022) to compress the pseudo videos into latent discrete tokens. C-ViViT is trained by reconstructing the pseudo videos with L2 reconstruction loss (Kingma & Welling, 2013), Vector Quantization (VQ) loss (Van Den Oord et al., 2017), GAN style adversarial loss (Karras et al., 2020), and image perceptual loss (Johnson et al., 2016; Zhang et al., 2018). For the latent space created by a C-ViViT, we consider two generative models to fit a prior $p(z)$ for the latent tokens:
@@
Network Architectures. [^1] For 1-frame models, we have tried using deeper architectures but observed no improvement in performance, which suggests that our pseudo video framework can help further improve the performance of generative models while simply increasing the model size becomes ineffective. We train VQVAE based generative models with a codebook size of 1024 for the discrete latent tokens. Specifically, we use C-ViViT as compression model (autoencoder) for pseudo videos. For a pseudo video with shape $(T,H,W,C)$, we compress it to discrete latent tokens with shape $(\frac{T}{2},\frac{H}{4},\frac{W}{4},C)$. For images with shape $(1,H,W,C)$, we use VQVAE to compress them to tokens with shape $(1,\frac{H}{4},\frac{W}{4},C)$. We then consider two video generative models, VideoGPT and Phenaki (and their image generative model counterparts, ImageGPT and MaskGit), to fit the prior over the latent tokens.
  • VQVAE/C-ViViT (reconstruction). We use a similar architecture as in Villegas et al. (2022), which has a 4-layer spatial Transformer and a 4-layer temporal Transformer, with a hidden dimension of 512. For 1-frame VQVAE model, we consider a 8-layer spatial Transformer.
@@
Results. Table 2 shows the last-frame FID of C-ViViT for reconstruction (we also report Peak Signal-to-noise Ratio (PSNR) for reconstruction in Appendix D) and that of VideoGPT and Phenaki for AR and masked generation on CIFAR10 and CelebA, respectively. The 1-frame results correspond to the performance of their image model counterparts (i.e., VQVAE for image reconstruction, and ImageGPT and MaskGit for AR and masked image generation, respectively). We observe that pseudo videos indeed help improve the training of the C-ViViT as the reconstruction quality of the last frame is significantly improved with a few more frames. We show reconstructed CIFAR10 and CelebA images from different C-ViViT models in Figures LABEL:fig:recon_cifar10 and LABEL:fig:recon_celeba, respectively. It can be seen that for both datasets, the reconstructed images trained with 1-frame models are over-smoothed and this issue is resolved by using pseudo videos which produce much sharper images. Quantitatively, reconstruction FID improves as more frames are used. Pseudo videos also improve the image generation performance compared to the 1-frame results. Interestingly, we see a diminishing return as we include more frames. For CelebA images, 18-frame pseudo videos help achieve the best image generation performance for both AR and latent masked generation, but 8-frame models achieve a comparable performance. For CIFAR10 images, 8-frame pseudo videos result in better image generation performance than 18-frame pseudo videos, which suggests the latent codes have a more complex prior distribution in order to reconstruct model pseudo videos with 18 frames well and therefore this prior is more difficult for VideoGPT or Phenaki to capture. In summary. while pseudo videos help improve the performance, the optimal number of frames may depend on the dataset and the augmentation strategy. This diminishing return is not a severe issue in practice since practitioners may prefer to improve the generation with just a few more pseudo frames to avoid introducing a high computational cost.
Table 2: Last-frame FID of images produced by C-ViViT (reconstruction), VideoGPT (AR generation) and Phenaki (latent masked generation) trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQVAE (reconstruction), ImageGPT (AR generation) and MaskGit (latent masked generation) trained on original CIFAR10 and CelebA images.
@@ 4.1 Is First-order Markov Chain the Optimal Choice for Creating Pseudo Videos?
Moreover, the non-optimality of creating pseudo video via first-order Markov chain becomes clear: the first-order Markov data augmentation implies that $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$ and consequently $\mathcal{L}_{2}^{*}=\mathcal{L}_{1}^{*}$. More specifically, for strict inequality in Eq 14, we need to avoid $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$, which is equivalent to avoiding the use of either first-order Markov chain $x_{T}\rightarrow x_{T-1}\rightarrow x_{T-2}$ or $x_{T}\leftarrow x_{T-1}\rightarrow x_{T-2}$. This analysis is informative for us to design better pseudo videos, for example through data augmentation with higher-order Markov chains. We formalize the above informal reasoning into Theorem 4.1 and provide the formal proof in Appendix B.
@@ 4.2 Experiments
  • First-order Markov. We add Gaussian noise recursively 3 or 7 times to create first-order Markov pseudo videos with a linear schedule (Ho et al., 2020) with $\beta$ ranging from 0.0001 to 0.05: $x_{T-t}=\sqrt{1-\beta_{t}}x_{T-t+1}+\sqrt{\beta_{t}}\epsilon$, $\epsilon\sim\mathcal{N}(0,I)$.
  • High-order Markov. While using the same noise schedule to create $x_{T-t}$, instead of adding Gaussian noise to $x_{T-t+1}$, we use a simple strategy to create high-order Markov pseudo videos by adding Gaussian noise to the mean of $\{x_{T-t+s}\}_{s=1}^{t}$: $x_{T-t}=\sqrt{1-\beta_{t}}[\frac{1}{t}\sum_{s=1}^{t}x_{T-t+s}]+\sqrt{\beta_{t}}\epsilon$, $\epsilon\sim\mathcal{N}(0,I)$.
@@ 6 Conclusion and Discussions
Appendix A Reproducibility Statement
Appendix B Proof of Theorem 4.1
@@ Appendix A Reproducibility Statement
Appendix C Hyperparameters
Table 6: Hyperparamters used for C-ViViT architecture and optimizer.
1-frame 8-frame 18-frame
Number of spatial layers 8 4 4
Number of temporal layers - 4 4
Embedding dimension 512 512 512
Hidden dimension 512 512 512
Number of heads 8 8 8
Learning rate 1e-4 1e-4 1e-4
Learning rate scheduler Cosine decay Cosine decay Cosine decay
Number of training steps 100k 100k 100k
Batch size 64 64 64
Table 7: Hyperparamters used for VideoGPT architecture and optimizer.
1-frame 8-frame 18-frame
Number of layers 8 8 8
Embedding dimension 144 144 144
Hidden dimension 144 144 144
Number of heads 4 4 4
Learning rate 1e-4 1e-4 1e-4
Learning rate scheduler Cosine decay Cosine decay Cosine decay
Number of training steps 100k 100k 100k
Batch size 64 64 64
Table 8: Hyperparamters used for Phenaki architecture and optimizer.
1-frame 8-frame 18-frame
Number of layers 6 6 6
Embedding dimension 512 512 512
Hidden dimension 512 512 512
Number of heads 8 8 8
Learning rate 1e-4 1e-4 1e-4
Learning rate scheduler Cosine decay Cosine decay Cosine decay
Number of training steps 200k 200k 200k
Batch size 64 64 64
Table 9: Hyperparamters used for UNet architecture and optimizer for Video diffusion.
1-frame 4-frame 8-frame
Number of downsampling/upsampling layers 4 4 4
Number of residual blocks 2 2 2
Base channel size 128 128 128
Number of diffusion steps per generating a frame 1000 1000 1000
Learning rate 1e-4 1e-4 1e-4
Number of training steps 500k 500k 500k
Batch size 32 32 32
Appendix D PSNR for Reconstruction
Table 10: PSNR of reconstruction for the last frame of pseudo videos produced by C-ViViT trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQ-VAE trained on original CIFAR10 and CelebA images.
CIFAR10 CelebA
1-frame 8-frame 18-frame 1-frame 8-frame 18-frame
Reconstruction 17.20 21.86 25.89 20.52 24.66 29.68